OpenStudy (anonymous):

Does anyone know how you would go about solving an equation with the same variable on both sides? Like... y=5/(y+2)

OpenStudy (anonymous):

5+2=7

OpenStudy (anonymous):

written differently it would be y=sqrt{5-2y}

OpenStudy (anonymous):

multiple both sides by (y + 2) then the left side will be y^2 + 2y = 5.

OpenStudy (anonymous):

isn't there a simpler way then just plugging numbers in

OpenStudy (anonymous):

Then use quadratic formula to solve

OpenStudy (anonymous):

is that the only way? i thought there was a different trick

OpenStudy (anonymous):

for simple equation if you can stare at it and spits out numbers then OK, but for complicated equations you can't plug in numbers in the first place. You CAN plug in numbers once you have a solution just to check if you got the right answer

OpenStudy (anonymous):

what topic are you studying now?

OpenStudy (anonymous):

so the equation should be y^2 +2y -5 = 0. Use the quadratic formula if studied this already. It's pretty straight forward. No tricks.

OpenStudy (anonymous):

the real equation I'm trying to figure out is y^3+x^3=3xy for y explicitly